# A Fast Parallel Approach for Neighborhood-based Link Prediction by Disregarding Large Hubs

### v1

@@ 5.3. Performance Analysis
In this section, we examine the phase split of the LHub approach to identify further optimization opportunities. To do this, we generate an observed graph $E^{O}$ with $0.1|E|$ unobserved edges $E^{U}$ for each dataset graph using random division, as detailed in Section 5.1.4. For each observed graph, we predict $0.1|E|$ links using the LHub approach and the similarity measures outlined in Section 3.1, utilizing the appropriate hub limit $L_{H}$ value, as per Section 4.3.
Figure 6 shows that LHub, spends a majority of its runtime, $63\%$ on average, in the scoring phase. This is particularly notable in social networks with a high average degree. However, a substantial amount of time is still spent on the merging phase, which involves combining predicted edges with top-$k$ scores from each thread into a global top-$k$ list of predicted edges. The sequential nature of the merging phase likely contributes to this runtime, and addressing this aspect is a potential focus for future work.
@@ 5.4. Strong Scaling
In the final analysis, we evaluate the strong scaling performance of the LHub approach, where we perform link prediction while disregarding large hubs. The assessment involves varying the number of threads from $1$ to $32$ in multiples of $2$ for each input graph. We measure the average time taken to predict $0.1|E|$ links using similarity measures defined in Section 3.1, incorporating the best hub limit $L_{H}$ setting identified in Section 4.3. Figure 6 presents the results, illustrating not only the overall scaling performance but also the scaling of the two phases of each link prediction method: identifying edges with top-$k$ scores in each thread (scoring phase) and combining scores across threads to obtain the global top-$k$ edges (merging phase).
With $32$ threads, the LHub approach achieves an overall speedup of 10.4×10.4$\times$10.4 × compared to sequential execution, indicating a performance increase of 1.6×1.6$\times$1.6 × for every doubling of threads. The scalability is limited, as the cost of the merging phase increases with an increase in the number of threads, and because the merging is performed with sequential execution. In fact, at $32$ threads, the merging phase obtains no speedup of 1.0×1.0$\times$1.0 ×, while the scoring phase achieves a speedup of 17.3×17.3$\times$17.3 ×.
@@ 6. Conclusion
In this report, we introduced the IHub approach, an enhanced parallel method that efficiently finds common neighbors and handles large graphs by tracking top-$k$ edges per-thread and later merging them globally. Additionally, we presented the LHub heuristic approach, which disregards large hubs. It is based on the notion that low-degree nodes contribute significant similarity among neighbors, in contrast to high-degree nodes (interestingly, this is similar to the idea behind the AA and RA similarity metrics). We then experimentally determined suitable hub limits for link prediction with each similarity metric.
Our results indicate that the LHub approach is, on average, over 963×963$\times$963 × and 163×163$\times$163 × faster than the IHub approach with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. It achieves this speedup while predicting links with an average F1 score that is $61\%$ higher and $16\%$ lower, respectively — meaning similar F1 scores without being too low or high. Notably, on the sk-2005 graph with $0.1|E|$ edges removed, LHub achieves a link prediction rate of $38.1M$ edges/s.
Moreover, RA metric shows superior performance in terms of F1 score and runtime on web graphs and social networks using the IHub approach. On road networks and protein k-mer graphs, JC similarity metric outperforms others with the IHub approach. With the LHub approach and $10^{-2}|E|$ unobserved edges, CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks, LHN metric (hub limit $L_{H}$ of $4$) performs best on road networks, and JC metric (hub limit $L_{H}$ of $256$) is optimal for protein k-mer graphs. However, with $0.1|E|$ unobserved edges, CN metric (hub limit $L_{H}$ of $32$) is the best across all graphs. LHub outperforms IHub, achieving higher F1 scores, particularly on web graphs and social networks.
When predicting $0.1|E|$ edges with the LHub approach, we observed that $63\%$ of the runtime is spent on the scoring phase, and especially higher on social networks, with high average degree. However, a significant portion is also dedicated to the merging phase, suggesting potential optimization opportunities for future work. In the strong scaling analysis, the LHub approach achieves a speedup of 10.4×10.4$\times$10.4 × with 32 threads, indicating a performance increase of 1.6×1.6$\times$1.6 × for every doubling of threads. However, scalability is constrained, particularly due to the merging phase.

### v3

@@ 5.3. Performance Analysis
In this section, we examine the phase split of the DLH approach to identify further optimization opportunities. To do this, we generate an observed graph $E^{O}$ with $0.1|E|$ unobserved edges $E^{U}$ for each dataset graph using random division, as detailed in Section 5.1.4. For each observed graph, we predict $0.1|E|$ links using the DLH approach and the similarity measures outlined in Section 3.1, utilizing the appropriate hub limit $L_{H}$ value, as per Section 4.3.
Figure 5 shows that DLH, spends a majority of its runtime, $63\%$ on average, in the scoring phase. This is particularly notable in social networks with a high average degree. However, a substantial amount of time is still spent on the merging phase, which involves combining predicted edges with top-$k$ scores from each thread into a global top-$k$ list of predicted edges. The sequential nature of the merging phase likely contributes to this runtime, and addressing this aspect is a potential focus for future work.
@@ 5.4. Strong Scaling
In the final analysis, we evaluate the strong scaling performance of the DLH approach, where we perform link prediction while disregarding large hubs. The assessment involves varying the number of threads from $1$ to $32$ in multiples of $2$ for each input graph. We measure the average time taken to predict $0.1|E|$ links using similarity measures defined in Section 3.1, incorporating the best hub limit $L_{H}$ setting identified in Section 4.3. Figure 6 presents the results, illustrating not only the overall scaling performance but also the scaling of the two phases of each link prediction method: identifying edges with top-$k$ scores in each thread (scoring phase) and combining scores across threads to obtain the global top-$k$ edges (merging phase).
With $32$ threads, the DLH approach achieves an overall speedup of 10.4×10.4$\times$10.4 × compared to sequential execution, indicating a performance increase of 1.6×1.6$\times$1.6 × for every doubling of threads. The scalability is limited, as the cost of the merging phase increases with an increase in the number of threads, and because the merging is performed with sequential execution. In fact, at $32$ threads, the merging phase obtains no speedup of 1.0×1.0$\times$1.0 ×, while the scoring phase achieves a speedup of 17.3×17.3$\times$17.3 ×.
@@ 6. Conclusion
In this technical report, we first improved upon the baseline approach (IBase). Our parallel IBase approach efficiently finds common neighbors and handles large graphs by tracking top-$k$ edges per-thread and later merging them globally. Additionally, we presented a novel heuristic approach, which disregards large hubs (DLH). It is based on the notion that low-degree nodes contribute significant similarity among neighbors, in contrast to high-degree nodes (interestingly, this is similar to the idea behind the AA and RA similarity metrics). We then experimentally determined suitable hub limits for link prediction with each similarity metric.
Our results indicate that the DLH approach is, on average, over 1622×1622$\times$1622 × and 415×415$\times$415 × faster than the IBase approach with $10^{-2}|E|$ and $0.1|E|$ unobserved edges, respectively. It achieves this speedup while predicting links with an average F1 score that is $80\%$ higher and $13\%$ lower, respectively — meaning similar F1 scores without being too low or high. Notably, on the sk-2005 graph with $0.1|E|$ edges removed, DLH achieves a link prediction rate of $38.1M$ edges/s.
Moreover, RA metric shows superior performance in terms of F1 score and runtime on web graphs and social networks using the IBase approach. On road networks and protein k-mer graphs, JC similarity metric outperforms others with the IBase approach. With the DLH approach and $10^{-2}|E|$ unobserved edges, CN metric (hub limit $L_{H}$ of $32$) excels on web graphs and social networks, LHN metric (hub limit $L_{H}$ of $4$) performs best on road networks, and JC metric (hub limit $L_{H}$ of $256$) is optimal for protein k-mer graphs. However, with $0.1|E|$ unobserved edges, CN metric (hub limit $L_{H}$ of $32$) is the best across all graphs. DLH outperforms IBase, achieving higher F1 scores, particularly on web graphs and social networks.
When predicting $0.1|E|$ edges with the DLH approach, we observed that $63\%$ of the runtime is spent on the scoring phase, and especially higher on social networks, with high average degree. However, a significant portion is also dedicated to the merging phase, suggesting potential optimization opportunities for future work. In the strong scaling analysis, the DLH approach achieves a speedup of 10.4×10.4$\times$10.4 × with 32 threads, indicating a performance increase of 1.6×1.6$\times$1.6 × for every doubling of threads. However, scalability is constrained, particularly due to the merging phase.